### v1

# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

Abstract
Real-time prediction of corridor-level crash occurrence risk (COR) remains challenging, as existing near-miss based extreme value models oversimplify collision geometry, exclude vehicle–infrastructure (V–I) interactions, and inadequately capture spatial heterogeneity in vehicle dynamics. This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a non-stationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors. High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme V–V and V–I near misses. The model incorporates dynamic variables and roadway features as covariates, with partial pooling across locations to address unobserved heterogeneity. Results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near-misses. Predictive validation using ROC–AUC confirms strong performance: 0.89 for V–V segments, 0.82 for V–V intersections, 0.79 for V–I segments, and 0.75 for V–I intersections. Model interpretation reveals that relative speed and distance dominate V–V risks at intersections and segments, with deceleration critical in segments, while V–I risks are driven by speed, boundary proximity, and steering/heading adjustments. These findings highlight the value of a statistically rigorous, geometry-sensitive, and spatially adaptive modeling approach for proactive corridor-level safety management, supporting real-time interventions and long-term design strategies aligned with Vision Zero.
keywords:
[…]
1 Introduction
The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches. Urban corridors are particularly vulnerable, where a single incident can disrupt mobility, trigger cascading congestion, and impose severe economic losses. In 2023, traffic crashes in the United States claimed 40,901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes is estimated to exceed \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite extensive research on crash patterns, risk estimation, and mitigation strategies (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b), accurate prediction of crash risk in urban corridors remains methodologically underdeveloped.
Most existing models focus exclusively on vehicle–vehicle (V–V) near-misses, largely due to the availability of vehicle trajectory data, while systematically excluding vehicle–infrastructure (V–I) interactions such as lane edges, curbs, barriers, and crosswalks. Yet V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021). This omission leaves corridor-level safety models inherently incomplete. Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
Crash risk prediction strategies based on historical retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011) suffer from well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time conditions. These challenges motivate a paradigm shift toward proactive, trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). Derived from high-resolution vehicle trajectory data, SSMs reflect underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among SSMs, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, these formulations often assume constant velocity and linear trajectories, which bias results in boundary-constrained environments. Recent research has extended TTC into higher-fidelity two-dimensional (2D) indicators (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration. Yet, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence near-miss likelihood and severity. To support reliable short-term crash occurrence risk (COR) estimation, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure. This study builds on these efforts by introducing a high-fidelity, geometry-aware 2D-TTC indicator that explicitly incorporates vehicle geometry, dynamic states, and roadway boundaries, ensuring realistic and efficient detection of both V–V and V–I near misses.
While SSMs capture frequent near-misses, a robust COR estimation framework requires extrapolating extreme near-misses to rare event outcomes. To date, several studies have proposed probabilistic (Saunier and Sayed, 2008), causal (Davis et al., 2011), and extreme value theory (EVT) (Fu and Sayed, 2021a, b; Zheng et al., 2021b; Ali et al., 2023a; Zheng et al., 2021a; Songchitruksa and Tarko, 2006; Zheng et al., 2014; Tarko, 2012). EVT provides a principled statistical foundation for this task, modeling the tail behavior of the extreme near-miss distribution. Two EVT approaches dominate: the Generalized Extreme Value (GEV) distribution using Block Maxima (BM) sampling and the Generalized Pareto Distribution (GPD) using Peaks Over Threshold (POT) sampling (Coles et al., 2001). The BM approach is particularly suitable for real-time analysis due to its short block structure, lower threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
Early EVT applications in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area. More recently, non-stationary GEV models have been introduced, allowing covariates to influence location and scale parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby capture heterogeneity across traffic states (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions have further addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). However, most applications remain geographically narrow and functionally limited to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024). Although some studies have incorporated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), these frameworks generally omit explicit integration of roadway geometry and directional heterogeneity at the corridor level. Such integration requires a modeling structure capable of jointly representing vehicle dynamics (which vary temporally) and roadway features (which vary spatially across segments and intersections). Therefore, a critical gap remains for a comprehensive Bayesian hierarchical GEV framework capable of jointly modeling vehicle dynamics, roadway geometry, flow directionality, and multilevel heterogeneity. Such a framework is essential for accurate corridor-level COR estimation that reflects the complex spatiotemporal dynamics of urban traffic environments. This paper makes the following unique contributions:
  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in prior studies.
  • Introduces a high-fidelity, geometry-aware 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
  • Advances a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) Univariate GEV model that integrates roadway geometry, traffic flow directionality, and multilevel heterogeneity, enabling corridor-wide yet location-sensitive crash risk estimation
Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for a proactive COR framework in urban corridors, leveraging Arogovers-2 AV trajectory and HD map data to support both operational interventions and long-term safety planning.
The remainder of this paper introduces the proposed 2D-TTC indicators, the HBSGRP–UGEV modeling framework, and the data preparation used in the analysis. This is followed by the presentation of model estimation, validation, and comparative results. The paper concludes with key findings, limitations, and directions for future research.
2 Near-miss detection framework
This section presents the modeling framework for computing a geometry-aware, high-fidelity 2D-TTC to support COR estimation. The framework integrates a continuous-time kinematic bicycle model, vehicle geometry representation, and a spatial near-miss detection algorithm, allowing capture of V-V and V-I interactions. These components collectively enable the accurate forecasting of vehicle trajectories and the identification of potential near misses. The resulting architecture forms the computational backbone of the SSM analysis, ensuring a realistic representation of near-miss events in mixed-traffic environments.
2.1 Vehicle dynamics model
The vehicle dynamics module provides the predictive foundation of the framework. This adopts the kinematic bicycle model to represent vehicle footprint, as it is suitable for short-term near-miss analysis, where steering and acceleration dominate, and lateral slip can be neglected. The motion is represented by a four-dimensional (4D) dynamic state vector. At any time $t$, the state of vehicle $i$ is given by:
$\mathbf{x}_{i}(t)=\begin{bmatrix}x_{i}(t)\\[2.84526pt] y_{i}(t)\\[2.84526pt] \theta_{i}(t)\\[2.84526pt] v_{i}(t)\end{bmatrix}$ (1)
Here, $x_{i}(t),y_{i}(t)$ denote the vehicle’s global center coordinates, $\theta_{i}(t)$ its heading angle, and $v_{i}(t)$ its speed. Over the prediction window, each vehicle is assumed to maintain constant acceleration $a_{i}$ and steering angle $\delta_{i}$, simplifying short-horizon trajectory integration. The motion then follows the standard kinematic bicycle equations:
$\dot{x}_{i}(t)$ $=v_{i}(t)\cos\theta_{i}(t)$ (2)
$\dot{y}_{i}(t)$ $=v_{i}(t)\sin\theta_{i}(t)$ (3)
$\dot{\theta}_{i}(t)$ $=\frac{v_{i}(t)}{L_{i}}\tan\delta_{i}$ (4)
$\dot{v}_{i}(t)$ $=a_{i}$ (5)
Here, $L_{i}$ is the vehicle wheelbase, influencing the turning radius and path curvature of vehicle $i$. The steering angle $\delta_{i}$ and acceleration $a_{i}$ are treated as fixed inputs over the lookahead period. Equations (2-5) describe the rates of change in position, orientation, and velocity, determining the vehicle’s trajectory. To model interactions, individual states are concatenated into an eight-dimensional (8D) joint state vector, defined as:
$X(t)=\begin{bmatrix}x_{A}(t)&y_{A}(t)&\theta_{A}(t)&v_{A}(t)&x_{B}(t)&y_{B}(t)&\theta_{B}(t)&v_{B}(t)\end{bmatrix}^{\top}$ (6)
The vector field that governs the evolution of this combined state is F{F} defined as:
$\dot{X}(t)=F\bigl{(}X(t)\bigr{)}=\begin{bmatrix}v_{A}\cos\theta_{A}&v_{A}\sin\theta_{A}&\frac{v_{A}}{L_{A}}\tan\delta_{A}&a_{A}&v_{B}\cos\theta_{B}&v_{B}\sin\theta_{B}&\frac{v_{B}}{L_{B}}\tan\delta_{B}&a_{B}\end{bmatrix}^{\top}$ (7)
The state-space vector field in Eqn.7 gives the instantaneous rate of change of the joint state $\dot{X}(t)$. Its forward evolution is computed using the fourth-order Runge–Kutta (RK4) method, which balances accuracy and efficiency for nonlinear vehicle dynamics. The time horizon is discretized into $N$ steps of size $\Delta t$ ($t_{n}=n\Delta t$, $n=0,1,\dots,N$), and the RK4 update for state ${X}(t_{n+1})$ from ${X}(t_{n})$ is:
$X_{n+1}=X_{n}+\frac{\Delta t}{6}\left(k_{1}+2k_{2}+2k_{3}+k_{4}\right)$ (8)
where the intermediate slope evaluations ${k}_{1},\dots,{k}_{4}\in\mathbb{R}^{8}$ are defined as:
$k_{1}$ $=F(X_{n})$
$k_{2}$ $=F\left(X_{n}+\frac{\Delta t}{2}k_{1}\right)$
$k_{3}$ $=F\left(X_{n}+\frac{\Delta t}{2}k_{2}\right)$
$k_{4}$ $=F\left(X_{n}+\Delta t\cdot k_{3}\right)$ (9)
This integration is repeated for each timestep $n\in\{0,\dots,N-1\}$, producing a sequence of predicted states $\{{X}_{n}\}$ over the horizon. These trajectories form the basis for geometric collision checks between vehicles and roadway boundaries.
2.2 Vehicle geometry representation
While the dynamic model predicts vehicle positions and headings, accurate near-miss detection requires a realistic representation of the vehicle footprint. Simplifying vehicles as points or circles often misrepresents interactions, especially during lane changes, swerving, or angled approaches. To address this, each vehicle is modeled as a rigid rectangle defined by its actual length $L_{i}^{\text{veh}}$ and width $W_{i}^{\text{veh}}$ (vehicle $i\in\{A,B\}$). The vehicle is centered at its center of gravity, with four corners specified in the body-fixed coordinate system as:
${b}_{i}^{(j)}=\begin{bmatrix}\xi_{i}^{(j)}\\ \eta_{i}^{(j)}\end{bmatrix},\quad j=1,2,3,4$ (10)
where $\xi_{i}^{(j)}\in\left\{+\frac{L_{i}^{\text{veh}}}{2},\ -\frac{L_{i}^{\text{veh}}}{2}\right\}$ and $\eta_{i}^{(j)}\in\left\{+\frac{W_{i}^{\text{veh}}}{2},\ -\frac{W_{i}^{\text{veh}}}{2}\right\}$. This results in four corners: front-left, front-right, rear-left, and rear-right, expressed relative to the vehicle’s center in the local frame.
To compute the position of each corner in the global coordinate frame at time $t$, the body-fixed corner coordinates are first rotated by the vehicle’s heading angle $\theta_{i}(t)$, and then translated by the vehicle’s center position $(x_{i}(t),y_{i}(t))$. This transformation is written as:
${p}_{i}^{(j)}(t)=\begin{bmatrix}x_{i}(t)\\ y_{i}(t)\end{bmatrix}+R(\theta_{i}(t))\cdot{b}_{i}^{(j)}$ (11)
where $R(\theta_{i}(t))$ is the 2D rotation matrix defined as:
$R(\theta_{i}(t))=\begin{bmatrix}\cos\theta_{i}(t)&-\sin\theta_{i}(t)\\ \sin\theta_{i}(t)&\cos\theta_{i}(t)\end{bmatrix}$ (12)
This transformation rotates the local corner coordinates by the heading angle and translates them into the global frame. Repeating it for all corners at each timestep produces the vehicle’s time-varying geometric footprint, which serves as input for near-miss detection. These footprints enable precise evaluation of proximity-based events, both V-V and V-I, ensuring accurate 2D-TTC estimation while reducing false positives from geometric simplifications.
2.3 V-V near-miss detection
Vehicles are modeled as rigid rectangles, and their interactions are determined by predicted states and orientations. At each timestep of the RK4 integration, the global positions of all four corners are computed using the kinematic and rotational transformations. A collision is defined to occur when a corner of one vehicle lies within a proximity threshold of a corner of another. Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$. A V–V near-miss is flagged if any pair of these corners satisfies the proximity condition.
${p}_{A}^{(j)}(t)={p}_{B}^{(k)}(t)$ (13)
An exact match between corners of two vehicles is highly unlikely due to floating-point arithmetic and continuous motion. Instead, a practical criterion is applied: a collision is flagged if the difference in both the x and y coordinates of any corner pair is less than or equal to a small threshold $\epsilon$:
$|x_{A}^{(j)}-x_{B}^{(k)}|$ $\leq\varepsilon$ (14)
$|y_{A}^{(j)}-y_{B}^{(k)}|$ $\leq\varepsilon$ (15)
The rotation–translation transformation from each vehicle’s center and heading angle gives the global corner coordinates, for vehicles $A$ and $B$:
$x_{A}^{(j)}$ $=x_{A}+\xi_{A}^{(j)}\cos\theta_{A}-\eta_{A}^{(j)}\sin\theta_{A}$ (16)
$y_{A}^{(j)}$ $=y_{A}+\xi_{A}^{(j)}\sin\theta_{A}+\eta_{A}^{(j)}\cos\theta_{A}$ (17)
$x_{B}^{(k)}$ $=x_{B}+\xi_{B}^{(k)}\cos\theta_{B}-\eta_{B}^{(k)}\sin\theta_{B}$ (18)
$y_{B}^{(k)}$ $=y_{B}+\xi_{B}^{(k)}\sin\theta_{B}+\eta_{B}^{(k)}\cos\theta_{B}$ (19)
At each timestep, all 16 corner-pair combinations are evaluated. A V–V near-miss is flagged if any pair satisfies the proximity condition, and the corresponding collision time $t_{c}$ is recorded for 2D-TTC computation. This discrete-time approach enables accurate detection of spatial overlap while avoiding the oversimplifications of circular vehicle approximations.
2.4 V-I near-miss detection
In addition to V-V interactions, the framework also detects near-misses between vehicles and roadway infrastructure such as curbs, medians, or barriers. Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
Road boundaries are represented as static polylines of $M$ discrete points $(X_{\ell},Y_{\ell})$ consistent with HD map formats in datasets such as Argoverse-2. At each timestep, the global positions of the four vehicle corners ${p}_{i}^{(j)}(t)$ are obtained from the predicted state. The distance between a vehicle corner $j$ and boundary point $\ell$, the distance at time $t$ is computed as:
$d_{j,\ell}(t)=\sqrt{(x^{(j)}(t)-X_{\ell})^{2}+(y^{(j)}(t)-Y_{\ell})^{2}}$ (20)
A near-miss is flagged if $d_{j,\ell}(t)\leq\varepsilon$ for any corner and boundary point, where is a proximity threshold accounting for numerical tolerances. When this occurs, the corresponding time $t_{c}$ is recorded as the instant of boundary contact. These events are then passed to the 2D-TTC framework to support extreme risk estimation.
2.5 Spatiotemporal COR
The near-miss detection framework integrates vehicle dynamics, geometric transformations, and spatial proximity checks to identify both V–V and V–I events. Forward trajectories are simulated using RK4 integration, vehicle corners are transformed into global coordinates, and geometric overlaps are evaluated to detect potential near-misses across the prediction horizon. The procedure is summarized in Algorithm 1, and a conceptual illustration is provided in Fig. 2.5. These steps yield a time series of 2D-TTC values for each interaction block.
Input: Initial state $\mathbf{X}(0)$, vehicle parameters $l_{i}$, $L_{i}^{\text{veh}},W_{i}^{\text{veh}}$, control inputs $a_{i}$, $\delta_{i}$, and boundary points $\{(X_{\ell},Y_{\ell})\}$
Output: Earliest collision time $t_{c}$
1
2for each timestep $n=0$ to N1N-1 do
3   Simulate $\mathbf{X}_{n}\rightarrow\mathbf{X}_{n+1}$ using RK4 integration;
4   Compute global corner positions $\mathbf{p}_{i}^{(j)}$ for each vehicle;
5
6 for each corner pair $(j,k)\in\{1,2,3,4\}^{2}$ do
7    if $|x_{A}^{(j)}-x_{B}^{(k)}|\leq\varepsilon$ or $|y_{A}^{(j)}-y_{B}^{(k)}|\leq\varepsilon$ then
8         Record V-V near-miss, set $t_{c}$;
9
10
11
12 for each corner $j$ and boundary point $\ell$ do
13      Compute $d_{j,\ell}(t)=\sqrt{(x^{(j)}-X_{\ell})^{2}+(y^{(j)}-Y_{\ell})^{2}}$;
14    if $d_{j,\ell}(t)\leq\varepsilon$ then
15         Record V-I near-miss, set $t_{c}$;
16
17
18
return the earliest collision time $t_{c}$ as 2D-TTC
Algorithm 1 Near-miss detection based on state integration and geometric overlap
Fig. 1.: Computation of 2D-TTC for near-miss scenarios
While this geometric framework enables precise localization of near-misses, it does not capture the probabilistic structure or temporal evolution of COR. To address this, the extracted blockwise minima of 2D-TTC are modeled using the EVT model. For each block, the model estimates the probability that 2D-TTC falls below a critical safety threshold, which defines the COR metric as a short-term indicator of crash risk. Block duration is chosen to balance temporal sensitivity with statistical stability across diverse contexts. This framework, therefore, combines deterministic near-miss detection with probabilistic EVT modeling to enable localized, data-driven assessment of crash-prone conditions. The complete methodological pipeline, from trajectory simulation to dynamic risk estimation, is summarized in Fig. 2.5.
Fig. 2.: COR time series generation
3 Extreme value theory (EVT)
Extreme Value Theory (EVT) provides the statistical foundation for estimating COR by extrapolating rare, high-severity events from frequent near-misses. In this study, extreme events are represented by block maxima of 2D-TTC derived from high-resolution vehicle trajectories (Algorithm 1). EVT allows risk estimation without requiring large volumes of historical crash data (Zheng and Sayed, 2020; Fu and Sayed, 2022a). Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD). While POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete, sensor-derived datasets. By contrast, BM aligns naturally with real-time AV data, avoids subjective thresholds, and supports scalable, time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006). Accordingly, this study adopts the BM–GEV framework (see Fig. 2.5).
The GEV has been widely applied in traffic safety research. Early studies modeled near-missets with univariate and multivariate GEV formulations (Orsini et al., 2020; Zhang et al., 2019; Zheng et al., 2014). Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b). Bayesian hierarchical models have further addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts. The increasing availability of high-resolution AV trajectory and HD map data has highlighted the need for flexible GEV formulations that jointly capture V–V and V–I interactions.
Despite this progress, most UGEV models still neglect the joint influence of roadway geometry and vehicle dynamics, two critical but distinct sources of heterogeneity. Roadway geometry (e.g., lane count, curvature, median type) is spatially structured and relatively stable within segments or intersections, while vehicle dynamics (e.g., speed, acceleration) vary temporally and spatially, driving short-term behavioral volatility. Capturing this multilevel heterogeneity is crucial for estimating corridor- and network-level crash risk. To address this gap, this study develops a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) UGEV model. Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segment types, corridor directions) while allowing for within-group variability. This structure provides both flexibility and parsimony, making it well-suited for high-resolution, corridor-scale crash risk modeling (Washington et al., 2020; Cai et al., 2018; Islam et al., 2023).
3.1 Univariate generalized extreme value (UGEV)
As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most safety-critical near-miss events in both V–V and V–I interactions. To implement this framework, continuous vehicle trajectories are partitioned into fixed-duration windows (e.g., 11 s). Within each window, the minimum 2D-TTC is extracted and negated, consistent with the upper-tail modeling convention of UGEV (Coles et al., 2001). The resulting sequence of BM serves as the input for parameter estimation. Throughout this study, we focus on maxima, denoted as $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d.) random variables, $X_{1},X_{2},\ldots,X_{n}$. To align with conventional UGEV formulations, we consider the negative values of the original variables derived from Algorithm 1, thereby defining the block maximum as $M_{n}=\max\{-X_{1},-X_{2},\ldots,-X_{n}\}$. Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-a_{n})/b_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function. According to EVT, the limiting distribution $G(x)$ must belong to one of three families: Gumbel, Fréchet, or Weibull, each of which can be captured by the parametric form of the GEV distribution (Coles et al., 2001). Let the UGEV be defined as:
$G(x;\mu_{x},\sigma_{x},\xi_{x})=\exp\left\{-\left[1+\xi_{x}\left(\frac{x-\mu_{x}}{\sigma_{x}}\right)\right]^{-1/\xi_{x}}\right\},\quad\text{where }1+\xi_{x}\left(\frac{x-\mu_{x}}{\sigma_{x}}\right)>0$ (21)
Here, $\mu_{x}\in\mathbb{R}$ is the location parameter, $\sigma_{x}>0$ the scale, and $\xi_{x}\in\mathbb{R}$ the shape. The value of $\xi_{x}$ determines tail heaviness: $\xi_{x}>0$ corresponds to the heavy-tailed Fréchet, $\xi_{x}<0$ to the bounded Weibull, and $\xi_{x}=0$ to the light-tailed Gumbel case.
3.1.1 Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model
To capture structured heterogeneity in extreme outcomes across the corridor, we employ a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model. Unlike conventional random-parameter models that assign variation to individual units, the HBSGRP formulation introduces variation at the group level. This structure allows systematic differences across predefined observational groups while maintaining parsimony and interpretability. In this specification, covariates are partitioned into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups. For a dataset divided into $K$ groups, with observation $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group. This enables the capture of localized relationships while preserving overall model coherence. The model is expressed within a three-layer hierarchical Bayesian framework. The data layer specifies the likelihood of observed block maxima of 2D-TTC from V–V and V–I interactions, assuming a UGEV distribution. Process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence. Prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
Bayesian inference proceeds by updating prior beliefs with observed data to obtain the posterior distribution. Formally, the joint posterior of parameters $\Theta$, given data $X$, is:
$q(\Theta\mid X)\propto q_{\text{data}}(X\mid\Theta)\cdot q_{\text{process}}(\Theta\mid\Psi)\cdot q_{\text{prior}}(\Psi)$ (22)
In this formulation, $q(\Theta\mid X)$ denotes the posterior distribution, integrating both the data likelihood and prior knowledge. The term $q_{\text{data}}(X\mid\Theta)$ corresponds to the likelihood function defined by the EVT assumptions. The component $q_{\text{process}}(\Theta\mid\Psi)$ captures the hierarchical dependency structure, including the group-varying parameters. Lastly, $q_{\text{prior}}(\Psi)$ encompasses the prior distributions of hyperparameters that control the model’s flexibility and penalize over-complexity.
The likelihood function for the data layer is specified as each extreme value sample $X_{k,i}$, observed for event $i$ within group $K$, is modeled as:
$q_{\text{data}}(X_{k,i}\mid\Theta)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sigma_{k,i}}\exp\left\{-\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1-1/\xi_{k,i}}$ (23)
The location ($\mu_{k,i}$), scale ($\sigma_{k,i}$), and shape ($\xi_{k,i}$) parameters are expressed as functions of group-level fixed-effect covariates $X_{k,j}$ (e.g., road geometry, speed limit) and sample-specific random-effect covariates $W_{k,i,j}$ (e.g., speed, acceleration, time-to-collision) as follows:
$\mu_{k,i}=\beta_{\mu,0}+\sum_{j=1}^{J}\beta_{\mu,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\mu,j,k}W_{k,i,j}$ (24)
$\vartheta_{k,i}=\beta_{\sigma,0}+\sum_{j=1}^{J}\beta_{\sigma,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\sigma,j,k}W_{k,i,j},\quad\sigma_{k,i}=\exp(\vartheta_{k,i})$ (25)
$\xi_{k,i}=\beta_{\xi,0}+\sum_{j=1}^{J}\beta_{\xi,j}X_{k,j}$ (26)
In the above equations $X_{k,j}$ defined as group-level fixed-effect covariate $j$, $W_{k,i,j}$is sample-specific random-effect covariate $j$. $\beta_{\mu,0},\beta_{\sigma,0},\beta_{\xi,0}$ as fixed intercepts for respective parameters. $\beta_{\mu,j},\beta_{\sigma,j},\beta_{\xi,j}$: fixed-effect coefficients. $\gamma_{\mu,j,k},\gamma_{\sigma,j,k}$: random-effect coefficients, varying across groups $k$. Therefore, the likelihood for the process layer, modeling within-group variability, is defined as:
$q_{\text{process}}(\Theta\mid\Psi)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sqrt{2\pi\tau_{\mu}^{2}}}\exp\left\{-\frac{1}{2\tau_{\mu}^{2}}(\mu_{k,i}-\bar{\mu}_{k})^{2}\right\}\\ \times\frac{1}{\sqrt{2\pi\tau_{\vartheta}^{2}}}\exp\left\{-\frac{1}{2\tau_{\vartheta}^{2}}(\vartheta_{k,i}-\bar{\vartheta}_{k})^{2}\right\}\times\frac{1}{\sqrt{2\pi\tau_{\xi}^{2}}}\exp\left\{-\frac{1}{2\tau_{\xi}^{2}}(\xi_{k,i}-\bar{\xi}_{k})^{2}\right\}$ (27)
where $\bar{\mu}_{k}$, $\bar{\vartheta}_{k}$, and $\bar{\xi}_{k}$ are group-specific average parameters.
Prior distributions are assigned to all model parameters. The priors for fixed and random effect coefficients are as follows:
$q_{\text{prior}}(\Psi)=q_{\beta_{\mu,0}}(\beta_{\mu,0})\cdot\prod_{j=1}^{J}q_{\beta_{\mu,j}}(\beta_{\mu,j})\cdot\prod_{k=1}^{K}\prod_{j=1}^{J^{\prime}}q_{\gamma_{\mu,j,k}}(\gamma_{\mu,j,k})\\ \times q_{\beta_{\sigma,0}}(\beta_{\sigma,0})\cdot\prod_{j=1}^{J}q_{\beta_{\sigma,j}}(\beta_{\sigma,j})\cdot\prod_{k=1}^{K}\prod_{j=1}^{J^{\prime}}q_{\gamma_{\sigma,j,k}}(\gamma_{\sigma,j,k})\ \times q_{\beta_{\xi,0}}(\beta_{\xi,0})\cdot\prod_{j=1}^{J}q_{\beta_{\xi,j}}(\beta_{\xi,j})$ (28)
Hyperpriors for the variance parameters $\tau_{\mu}^{2},\tau_{\vartheta}^{2},\tau_{\xi}^{2}$ are modeled using inverse gamma distributions:
$\tau_{\mu}^{2}\sim\text{IG}(\alpha_{\mu},\beta_{\mu}),\quad\tau_{\vartheta}^{2}\sim\text{IG}(\alpha_{\vartheta},\beta_{\vartheta}),\quad\tau_{\xi}^{2}\sim\text{IG}(\alpha_{\xi},\beta_{\xi})$ (29)
where $\text{IG}(\alpha,\beta)$ denotes an inverse gamma prior distribution ensuring positivity constraints for variance parameters.
3.2 COR estimation
Building on the HBSGRP–UGEV estimates, the framework compute COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block. Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, enabling a holistic assessment of potential COR under mixed traffic conditions.
Let $i$ denote block $i$ within group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments). The minimum 2D-TTC in each block is modeled with the HBSGRP-UGEV distribution, parameterized by $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$. Let $\omega$ be the critical 2D-TTC threshold below which a near-miss is considered crash-prone. The probability of a crash-level event in block $i$ of group $k$ is:
$P_{k,i}^{\text{crash}}=\Pr(2D-TTC_{k,i}\leq\omega)=1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\},\quad\text{where }1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)>0$ (30)
Assuming a homogeneous Poisson process for exceedances, the expected COR for group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$. Here $y_{k,i}$ is the number of near-misses (V-V or V-I) in block $i$, and $N_{k}$ the number of blocks in group $k$. The expected COR for group $k$ is:
$COR_{k}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot P_{k,i}^{\text{crash}}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (31)
The total expected crash frequency across the study area is then:
$CF_{\text{total}}=\sum_{k=1}^{K}COR_{k}=\frac{1}{T}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1+\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (32)
This formulation integrates exceedance probabilities from the HBSGRP–UGEV model with observed V–V and V–I near-miss counts, maintaining a direct link between real-world interaction intensity and modeled crash propensity. By separating probability from exposure, the framework yields interpretable, proactive crash frequency estimates aligned with EVT-based safety assessment.
4 Data description
The effectiveness of real-time COR estimation depends critically on the fidelity and coverage of trajectory data. Traditional sensing technologies, such as roadside cameras, loop detectors, and fixed LiDAR, have supported safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020). However, these systems suffer from narrow fields of view, occlusion, overlapping detections, and high maintenance costs, which limit scalability and data quality (St-Aubin et al., 2013). Recent advances in AV sensing offer a promising alternative. AV datasets offer high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers. This richness allows for realistic near-miss detection in complex traffic environments. Despite this potential, AV data remain underutilized in generalized frameworks for real-time COR estimation. For example, studies using the Waymo Open Dataset have largely focused on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
To address these challenges, this study utilizes the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and Washington, D.C.) using SAE Level 4 AVs. Argoverse-2 comprises more than 250k driving scenarios spanning 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared with Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions such as intersection near-misses, lane changes, and pedestrian crossings—conditions most indicative of crash risk. A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic. Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors. The sensor suite includes seven ring cameras and two front-facing stereo cameras (20 Hz) supported by dual 64-beam VLP-32C LiDAR sensors, producing  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
Table 1: Open-source AV datasets
Dataset Sensora Tasksb Hzc HD Map Scope / Notes
Argoverse-2 (Wilson et al. 2023) C, L D, T, S, M 10 Yes 250k scenarios across 113 segments; 11s each; 10 classes; 6 cities
nuScenes (Caesar et al. 2020) C, L, R D, T, S, M 2 No 40k scenarios across  1k segments; 23 classes; 2 cities
Waymo (Ettinger et al. 2021) C, L D, T, M 10 Yes 390k scenarios over  1950 segments; 12.6M trajectories; 6 cities
Lyft Level 5 (Kesten et al. 2019) C, L, R M 10 Yes 170k AV instances (25s) in one city; 10 classes
KITTI (Geiger et al. 2013) L D, T, M 10 No 15k scenes; classic full-stack AV benchmark
DAIR-V2X (Yu et al. 2022) L D Yes 71254 image frames for V2X scenes
OPV2V (Xu et al. 2022) C, L D, T, M 10 No 33k V2V simulation samples; 18k V2X frames (CARLA)
DeepAccident (Wang et al. 2024) C, L, R D, T, S, M 10 Yes 285k simulation samples incl. 57k V2X scenarios
Shift (Sun et al., 2022) C, L S, M 10 No 2.5M frames; 4850 segments; 8 cities
  • a
    Sensor: Camera (C), LiDAR (L), Radar (R).
  • b
    Tasks: Detection (D), Tracking (T), Segmentation (S), Motion Forecasting (M).
  • c
    Hz: Sampling frequency; “–” not reported.
Argoverse-2 was selected after benchmarking open-source AV datasets (Table 1). Waymo Open offers a greater total mileage but lower near-miss density and no AV-behavior labels. Lyft Level-5 is confined to a single city and lacks scene-level filtering. nuScenes samples at 2 Hz, limiting the capture of rapid interactions. In contrast, Argoverse-2 balances spatial coverage, sensor fidelity, near-miss-rich scenarios, and behavioral annotations, making it well-suited to this study’s goals. This study used motion-forecasting and HD-map modules of Argoverse-2 to extract vehicle trajectories, compute near-miss indicators, and encode roadway geometry.
4.1 Data Preprocessing
To operationalize the proposed framework, both the motion forecasting and HD map modules of the Argoverse-2 dataset were processed through a structured pipeline. The preprocessing consisted of four key stages: (i) extraction of multi-agent trajectories, (ii) transformation of coordinates from local map space to global geographic space, (iii) integration with HD map features to contextualize V-I interactions, and (iv) selection of a high-risk urban corridor for detailed analysis.
4.1.1 Motion forecasting dataset
Each motion forecasting scenario follows a hierarchical structure with three levels: scenario metadata, agent tracks, and frame-level states. Scenarios are 11s duration and are sampled at 10 Hz, producing 110 time steps per vehicle for fine-grained interaction modeling. At the scenario level, metadata include a unique scenario ID, the recording city, and the focal agent, typically the ego AV. Each scenario then contains multiple agent tracks, where each track corresponds to a road user and specifies the object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal, scored, or unscored), and a time-ordered sequence of states. Frame-level states capture position, velocity, and heading at each timestep, forming the basis for computing 2D-TTC.
To implement the framework, we developed a Python-based batch pipeline using the Argoverse-2 SDK. The pipeline parsed motion forecasting scenarios across all cities and extracted tracks for all agents—not just the focal vehicle—to retain critical multi-agent interactions. For each track, frame-level kinematic attributes were retrieved, including $(x,y)$ position, heading, velocity components, and timestamps.
Because positions were reported in local, city-specific frames, all coordinates were transformed to globally referenced WGS84. This involved (i) projecting each city’s local origin into its UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), and (ii) offsetting local $(x,y)$ positions and inverse-projecting them to latitude–longitude using pyproj. This geospatial normalization provided consistent referencing across cities and ensured compatibility with HD map overlays. Batch scripts automatically traversed scenario folders, performed attribute extraction and coordinate conversion, and exported unified CSV datasets containing both original city-frame and converted lat/lon values. These datasets were then used for spatial filtering, interaction detection, and overlay validation.
The processed files formed the basis for V–V and V–I near-miss detection. For each valid scenario, pairwise distances and relative approach rates were computed for all vehicle pairs, while V-I distances were derived from buffered polygons of lane edges and curbs in the HD map. Scenarios with no substantive interaction (e.g., straight-lane cruising without lane changes or turns) were excluded, ensuring the dataset emphasized behaviorally informative, high-risk conditions consistent with proactive safety modeling. Non-vehicle agents (e.g., pedestrians, cyclists) were also excluded at this stage, though the pipeline is extensible to incorporate vulnerable road users in future work.
4.1.2 HD map
The Argoverse-2 HD map dataset provides vectorized representations of roadway geometry and drivable space, enabling infrastructure-aware risk estimation. These maps are distributed as JSON files in local city coordinates and contain lane centerlines, left/right boundaries, crosswalk polygons, and drivable area definitions.
Lane-level geometry was extracted using the official Argoverse-2 API. The get_all_lane_segments() method returns centerline polylines along with left/right boundaries that define the navigable road space. Additional features, such as crosswalk polygons and drivable area masks, were also retrieved to support validation and visualization, though only lane boundaries and centerlines were directly used in near-miss detection.
Because all coordinates are expressed in city-specific Cartesian frames, global georeferencing was required to align HD maps with trajectory data. For each city, the WGS84 geodetic origin was projected into its respective UTM zone (e.g., Zone 17 for Miami and Pittsburgh, Zone 14 for Austin, Zone 10 for Palo Alto, Zone 18 for Washington, D.C.) using the pyproj library. Each local $(x,y)$ coordinate was then offset by this origin and inverse-projected into latitude–longitude, ensuring spatial consistency across cities. Once transformed, all HD map components: lane boundaries, centerlines, and crosswalks—were indexed using their global coordinates and converted into shapefiles or GeoPandas geometries. This geospatial transformation enabled efficient spatial querying, collision proximity analysis, and visual overlay with satellite imagery and trajectory data. In particular, the lane boundaries were instrumental in detecting V–I near-misses by allowing for the computation of lateral vehicle encroachments beyond the legal driving space. From the processed dataset, a high-density urban corridor in Miami was present to showcase the transformation (see Fig. 4.1.2). This illustrates the transformation of AV data into real-world coordinates, its alignment with HD maps and satellite imagery, and the resulting traffic patterns that highlight interaction hotspots across intersections and midblocks.
Fig. 3.: AV data transformed to real-world coordinates and aligned: (a) trajectories with lane boundaries and crosswalks, (b) validation with satellite imagery, and (c–f) directional traffic patterns highlighting interaction hotspots.
4.1.3 Trajectory preprocessing and derivation of vehicle dynamics
Before computing 2D-TTC, a series of preprocessing steps was applied to ensure the precision, continuity, and physical plausibility of vehicle motion data from the Argoverse-2 dataset. These steps included the extraction of raw trajectory data, trajectory smoothing and filtering, and the derivation of dynamic variables such as speed, acceleration, and steering angle. First, all multi-agent trajectories were transformed from local city-specific frames to global WGS84 coordinates, enabling consistent alignment with HD map geometries. Raw trajectories were then examined for noise caused by sensor jitter, occlusion, or missed detections—issues more common in human-driven vehicles than in AV tracks. To suppress such noise while preserving curvature and lane-change dynamics, cubic B-spline smoothing was applied to the position coordinates (Eilers and Marx, 1996; Choi et al., 2024). This method effectively removed local fluctuations while retaining key geometric features (see Fig.4.1.3). Stationary or irrelevant agents, such as parked vehicles and bicycles, were filtered out to reduce computational load and focus only on interaction-relevant entities.
Fig. 4.: Effect of B-spline smoothing on a sample trajectory, showing removal of noise while preserving geometric fidelity
Dynamic control inputs were then computed. Instantaneous speed $v_{t}$ was derived as the Euclidean norm of velocity components:
$v_{t}=\sqrt{v_{x,t}^{2}+v_{y,t}^{2}}$ (33)
Residual noise in $v_{t}$ was suppressed with a Savitzky–Golay filter (Chen et al., 2004) using a second-order polynomial and a 110-frame window (11s). Longitudinal acceleration ($a_{t}$) was obtained by first-order numerical differentiation:
$a_{t}=\frac{v_{t+1}-v_{t}}{\Delta t}$ (34)
where $\Delta t=0.01$s reflects the 10 Hz sampling rate. Heading $\theta_{t}$ was computed from directional vectors of each track and was unwrapped to maintain continuity across discontinuities at $\pm\pi$. The yaw rate $\dot{\delta}_{t}$ was derived from the kinematic bicycle model:
$\dot{\theta}_{t}=\frac{\theta_{t+1}-\theta_{t}}{\Delta t}$ (35)
$\delta_{t}=\tan^{-1}\left(\frac{L\cdot\dot{\theta}_{t}}{v_{t}}\right)$ (36)
The final dataset included per-frame values of speed, acceleration, heading, yaw rate, and steering angle for each track, providing stable and interpretable control inputs. These preprocessed trajectories formed the basis for subsequent near-miss detection using Algorithm 1 and the HBSGRP-UGEV model.
4.2 Study area
Under realistic and heterogeneous urban traffic conditions, a representative high-volume corridor was selected along Biscayne Boulevard in Miami, Florida. This arterial extends from Downtown Miami through the Upper East Side toward the northern city limits, serving as a critical connector to major activity centers, including the Miami International Airport region. It is characterized by multimodal interactions, geometric complexity, frequent turning maneuvers, and diverse traffic control features. The corridor was chosen for its rich coverage in the Argoverse-2 motion forecasting dataset, HD map, and its suitability for extracting sufficient samples for extreme near-miss scenarios.
As shown in Fig. 4.2, the study area was extracted using ArcGIS Pro based on scenario density, HD map fidelity, and heterogeneity of vehicle maneuvers. The extracted network includes a continuous corridor with multiple intersections, midblock segments, and bidirectional traffic operations. For analytical consistency, the corridor was subdivided into nine directional segments (S1–S9) and ten intersections (I1–I10). Segment boundaries were defined midblock to maintain flow continuity, while intersections were delineated to capture near-miss-prone zones such as turning paths, crosswalks, and lane merges. Each subregion was mapped with lane boundaries, crosswalks, and vehicle trajectories to preserve geometric fidelity and directional alignment. Insets in Fig. 4.2 illustrate the spatial configuration of the segmented units, which serve as the basis for grouped modeling and localized risk estimation.
Fig. 5.: Segmented study corridor along Biscayne Boulevard, Miami, with lane boundaries and crosswalks. Insets show direction-specific segments (S1–S9) and intersections (I1–I10)
Table 2 provides a detailed profile of the 28 study subregions, consisting of 10 intersections and 18 directional segments. Intersections generally show fewer scenarios and unique vehicles (e.g., I2 with 47 scenarios and 220 vehicles, I4 with 96 scenarios and 568 vehicles) compared to adjacent midblock segments, reflecting shorter dwell times but higher near-miss intensity. In contrast, segments accumulate a significantly higher number of scenarios and vehicles due to sustained travel. For example, S7 (NE 19-17 St) records 308 northbound and 332 southbound scenarios involving more than 4000 vehicles, while S8 (NE 17–15 St) captures nearly 4400 vehicles in 608 scenarios.
Table 2: Profile of study locations along Biscayne Boulevard
ID Location Direction Scenarios Unique vehicles Lane no Lane width (m) Driveway/Minor intersection Median type
Segment
S1 (NE 36–33 St) NB / SB 131 / 111 1706 / 813 3 / 3 3.5 3 Raised
S2 (NE 33–29 St) NB / SB 151 / 155 1532 / 1148 3 / 3 3.64 3 Raised
S3 (NE 29–26 St) NB / SB 191 / 186 1721 / 1295 3 / 3 3.36 4 Flush paved
S4 (NE 26–22 St) NB / SB 176 / 199 1542 / 1581 3 / 3 3.51 4 Flush paved
S5 (NE 22–21 St) NB / SB 127 / 123 705 / 643 3 / 2 3.3 No Flush paved
S6 (NE 21–19 St) NB / SB 235 / 245 1588 / 1861 3 / 3 3.62 1 Flush paved
S7 (NE 19–17 St) NB / SB 308 / 332 1704 / 2410 2 / 2 3.12 1 Flush paved
S8 (NE 17–15 St) NB / SB 294 / 314 1955 / 2406 3 / 3 3.57 2 Flush paved
S9 (NE 15–13 St) NB / SB 232 / 216 1565 / 1217 3 / 2 3.43 1 Flush paved
Intersection
I1 NE 36 St 97 743
I2 NE 33 St 47 220
I3 NE 29 St 132 699
I4 NE 26 St 96 568
I5 NE 22 St 106 640
I6 NE 21 St 134 647
I7 NE 19 St 214 1332
I8 NE 17 St 205 1328
I9 NE 15 St 252 1332
I10 NE 13 St 232 1139
Lane configurations vary between 2 and 3 lanes per direction, with average widths ranging from 3.1m (S7) to 3.6m (S6), indicating subtle but important differences in cross-sectional design. Driveway and minor intersection density also differs across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses. Median type alternates between raised (e.g., S1–S2) and flush paved (e.g., S3–S9), influencing turning maneuvers and midblock encroachments. Overall, the table highlights that intersections serve as near-miss-prone nodes with concentrated but lower-volume interactions, whereas segments provide richer samples of continuous traffic flow, characterized by greater geometric variability.
Building on the spatial segmentation and detailed subregion profiles outlined above, the next step involves quantifying dynamic safety conditions using the proposed 2D-TTC framework (Algorithm 1). For each block, extreme near-miss events were identified by forward simulating vehicle trajectories under the geometric and traffic conditions described in Table 2. Both V–V and V–I interactions were modeled to reflect the heterogeneous near-miss environments of intersections and segments. Trajectories were projected over a 3s horizon, and spatial overlaps with surrounding traffic or roadway boundaries were detected as potential near-miss events. Representative outcomes from this process are shown in Fig. 4.2, illustrating the spatial dynamics of predicted V–V and V–I near-misses within the Miami study corridor. These event visualizations provide the empirical basis for block-level 2D-TTC extraction across the 10 intersections (I1–I10), enabling localized evaluation of crash-prone conditions.
Table 3: Summary statistics for V–V and V–I near-misses at Intersection
Metric V–V V–I
Avg Min Max Std Avg Min Max Std
Continuous variables
TTC 1.617 0.11 2.99 0.863 1.304 0.10 2.99 0.889
Relative speed 5.257 0.00 23.724 4.446 8.753 0.00 22.276 4.035
Relative acceleration -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
Relative distance 11.673 2.419 46.857 6.844 1.520 0.60 3.80 1.110
Jerk -5.550 -13.609 9.315 2.652 -7.900 -10.800 2.800 14.700
Heading difference 1.077 0.00 6.251 1.410 -0.045 -3.138 3.141 1.647
Steering difference 0.409 0.00 3.142 0.666 -0.006 -1.571 1.571 0.130
Volume 13.462 2.00 31.00 5.969 10.461 1.00 31.00 5.392
Vehicle length 4.78 3.18 6.92 0.49 5.1 3.95 12.65 1.30
Vehicle width 2.22 1.76 3.16 0.30 2.04 1.36 3.08 0.23
Categorical variable: Near-miss location (Proportion %)
Through 72.29 0.00
Left Turn 25.99 46.08
Right Turn 8.90 53.92
(a) V-V
(b) V-I
Fig. 6.: Representative intersection-level near-miss detection using the 2D-TTC framework: (a) V–V and (b) V–I
Descriptive statistics of intersection-level V–V and V–I near-misses are summarized in Table 3. The continuous indicators include 2D-TTC, relative speed, acceleration, distance, jerk, and angular deviations (heading and steering), along with vehicle dimensions and exposure (volume). On average, 2D-TTC values for both near-miss types range between 1.3 and 1.6 s, with extreme cases falling below 0.1 s, reflecting highly critical near-miss conditions. V–I near-misses are associated with higher relative speeds and tighter distance gaps, indicating more abrupt encroachment against roadway boundaries. Conversely, V–V near-misses exhibit larger heading and steering deviations, along with higher jerk variability, suggesting unstable maneuvers such as sudden turns or evasive swerves. The categorical distribution of near-misses further illustrates key behavioral differences. Nearly three-quarters (72%) of V–V events occurred during through movements, with fewer during left (26%) or right (9%) turns. In contrast, V–I near-misses were almost entirely turning-related, with 46% occurring in left turns and 54% in right turns, and none during through movements. This pattern highlights that infrastructure-related near-misses typically occur during lateral maneuvers at intersections, where vehicles cross lane boundaries or encroach on curbs and medians.
Beyond intersections, the analysis was extended to directional roadway segments (S1–S9) using the same 2D-TTC framework (See Fig. 4.2).Table 4 summarizes descriptive statistics of both V–V and V–I near misses at the segment level. For V–V near-misses, average TTC values were longer (1.83 s) than those observed at intersections, suggesting greater temporal headway between vehicles in segment settings. However, the extreme minima (0.11 s) and high maximum relative speeds (up to 33.9 m/s) highlight the presence of aggressive overtaking and high-speed convergence. Relative distances also exhibited a much broader range (up to 141.5 m), indicating that near-misses often arise in more dispersed longitudinal interactions rather than tightly constrained intersection maneuvers.
Table 4: Summary statistics for V–V and V–I near-misses at segments
Metric V–V V–I
Avg Min Max Std Avg Min Max Std
Continuous variables
TTC 1.829 0.11 2.99 0.805 1.272 0.10 2.99 0.912
Relative speed 4.513 0.00 33.909 4.278 10.262 0.000 28.067 4.212
Relative acceleration -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
Relative distance 15.361 1.931 141.548 12.773 1.720 0.32 6.56 1.02
Jerk -7.121 -17.502 8.423 1.552 -5.70 -12.72 3.556 4.77
Heading difference 0.368 0.000 6.256 0.990 -0.032 -3.102 3.138 1.586
Steering difference 0.577 0.000 3.142 0.756 -0.001 -1.571 1.571 0.167
Volume 16.801 1.000 65.000 11.009 12.816 1.000 65.000 7.906
Vehicle Length 4.74 3.0 7.85 1.11 4.79 2.95 8.65 0.87
Vehicle width 2.12 1.36 3.08 0.23 2.22 1.76 4 0.18
Lane number 2.73 2 4 0.578 2.51 2.0 4.0 0.52
Lane width 3.28 2.99 3.96 0.14 3.27 2.99 3.96 0.141
Driveway intensity 0.009 0.000 0.015 0.005 0.009 0.000 0.015 0.005
Categorical variable: (Proportion %)
Through 58.55 0.00
Collision location Left Turn 9.52 46.08
Right Turn 31.93 53.92
Median type Divided 23.60 13.98
Undivided 76.4 86.02
(a) V-V
(b) V-I
Fig. 7.: Representative segment-level near-miss detection using the 2D-TTC framework: (a) V–V and (b) V–I
For V–I near-misses, the mean TTC was shorter (1.27s), with relative speeds peaking at 28.1 m/s, underscoring the abrupt nature of vehicle–boundary encounters. Compared with intersections, segment-based V–I near-misses involved longer approach distances but also showed greater instability in jerk and heading, reflecting frequent lateral drifts toward lane edges or curbs. These events often occurred near driveways and minor access points, where vehicles deviated from their intended path, creating encroachment risks. The categorical breakdown reinforces these patterns. For V–V events, nearly 59% of near-misses were associated with through movements, while about 32% and 10% were linked to right and left turns, respectively. In contrast, V–I near-misses were dominated by turning maneuvers, with more than half (53.9%) linked to right turns and 46% to left turns, and no near-misses detected during through movements. This distribution highlights the heightened role of lateral maneuvers in producing infrastructure-related risks. Furthermore, median type also influenced outcomes: 76% of V–V near-misses and 86% of V–I near-misses occurred along undivided sections, underscoring the vulnerability of open medians and flush-paved designs in segment environments.
Taken together, these segment-level findings reveal that while intersections are hotspots for tightly constrained V-V near-misses, segments carry distinctive risks tied to high-speed encounters and infrastructure encroachments. The observed distributions underscore the importance of treating intersections and segments as functionally distinct safety domains when modeling near-miss dynamics and designing targeted countermeasures.
(a) Intersection
(b) Segment: NB
(c) Segment: SB
Fig. 8.: BM near-miss frequencies by location type and direction
After identifying and compiling the BM near-miss events for each observation unit (segments and intersections), the frequencies of extreme V–V and V–I events were systematically aggregated and visualized in Fig. 4.2. This plot illustrates the spatial distribution of extreme near-miss frequencies across 28 locations, comprising 10 intersections and 18 segments, organized into intersection zones, primary-direction segments, and opposite-direction segments. This spatial aggregation confirms the integrity of the BM dataset and provides the foundation for modeling. The study extends prior approaches by incorporating both V–V and V–I near misses into the HBSGRP–UGEV framework, explicitly integrating roadway geometry and infrastructure features alongside vehicle dynamics. Prior research has emphasized the role of vehicle kinematics in crash risk (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015); here, additional geometric covariates, such as lane configuration and boundary proximity, are introduced. For example, high steering angles, often indicating abrupt lateral maneuvers—were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021). Fig. 4.2 highlights several notable spatial patterns. Opposite-direction segments (S82, S72, S62) exhibited the highest V–I frequencies, suggesting greater lateral variability and exposure to curb or median near-misses. Primary-direction segments (S11, S21, S81) showed elevated V–V near-miss frequencies, consistent with midblock environments dominated by longitudinal interactions such as close-following, lane changes, or overtaking. Among intersections, I7, I8, and I9 emerged as critical hotspots, with elevated frequencies of both near-miss types. This dual concentration reflects the operational complexity of intersections, where near-missing maneuvers, pedestrian flows, and signal control interact to amplify crash potential.
The distribution and magnitude of near-miss counts across all locations validate the effectiveness of the BM approach in capturing rare but safety-critical events. Specifically, nearly all study sites meet or exceed the recommended minimum threshold of 30 BM events for reliable GEV-based inference (Zheng et al., 2014), with the exception of intersection I2. This ensures that the dataset is statistically robust for subsequent HBSGRP–UGEV analysis. The spatial heterogeneity revealed in Fig. 4.2 further underscores the need to incorporate direction-specific geometric and behavioral covariates within the modeling framework. These results provide compelling justification for developing localized, high-resolution safety models that integrate both roadway configuration and observed vehicle dynamics, thereby enabling precise and context-sensitive real-time risk assessments in complex urban environments.
5 Modeling estimation result
This study develops a series of Bayesian hierarchical UGEV models to estimate COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami. Extreme near-misses are identified using the BM approach, wherein the minimum 2D-TTC within each temporal block is treated as the most critical interaction. The theoretical foundations and implementation of this sampling methodology are detailed in prior research (2025b). Each V–V and V–I interaction is treated as an independent block. Because extreme events are sparse at individual sites, direct location-specific estimation can yield unstable posteriors with wide uncertainty intervals. To address this, a grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while still preserving local heterogeneity. This clustering approach improves estimation stability and aligns with prior network-level applications of UGEV models in traffic safety (Zheng et al., 2014; Fu and Sayed, 2021c).
The primary analytical framework is the HBSGRP–UGEV model, while a simpler fixed-effect hierarchical model (HBSFP) was estimated as a benchmark. Accordingly, two models were developed for each near-miss type: (1) HBSFP and (2) HBSGRP. The grouped random parameter specification allows selected coefficients to vary across spatial clusters, thereby capturing latent differences linked to roadway geometry and interaction dynamics. Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), where lower values indicate stronger predictive performance (Zheng et al., 2019).
Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference. Its support for block updating and parallel execution makes it well-suited for capturing spatial dependence, random effects, and extreme value behavior within complex data structures. Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation, two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference. Convergence was verified both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, indicating satisfactory convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009). This estimation procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
The study incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the non-stationary UGEV distribution to reflect contextual and behavioral variability in COR. For both V–V and V–I interactions, dynamic covariates included relative speed, distance, acceleration, deceleration, headway difference, steering angle difference, jerk and traffic volume, factors capturing key dimensions of vehicle dynamics and near-miss severity. Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position for segment models; and a categorical indicator of turning movement for intersection models.
All covariates were specified in both $\mu$ and $\sigma$ components, enabling the models to capture variation in both the central tendency and dispersion of extreme 2D-TTC. In contrast, the shape parameter ($\xi$) was held constant across all locations. This choice is deliberate: $\xi$ governs tail heaviness in the UGEV distribution but is notoriously difficult to estimate from sparse BM samples. Allowing $\xi$ to vary with covariates often induces instability and overfitting (Coles et al., 2001; Cooley et al., 2006). Fixing $\xi$ improves convergence and yields more interpretable posterior distributions.
5.1 Posterior results and model performance
Tables 5 through 8 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-misses at intersections and segments. Across all cases, the HBSGRP specification consistently outperformed the HBSFP baseline, confirming its ability to capture unobserved heterogeneity and locatione-level variability. The DIC values for the HBSGRP models were 508.9 (V–V) and 3179.6 (V–I) at intersections, and 4189.0 (V–V) and 6550.0 (V–I) at segments. These correspond to relative reductions in DIC of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts. Such reductions exceed the widely accepted threshold of five DIC units, indicating statistically meaningful improvements in fit (El-Basyouny and Sayed, 2012b, a). To further assess model robustness, WAIC and LOOIC were also computed. Both measures produced rankings consistent with DIC, reinforcing the superior performance of the HBSGRP structure. By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion of near-miss severity, which fixed-effect structures fail to represent. Collectively, these results demonstrate that grouped random parameter models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
5.1.1 Intersection-level
The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions arising from V-V interactions at signalized locations. In the model specification stage, several covariates were tested as random effects; however, most roadway geometry variables exhibited negligible variation across sites and did not improve posterior fit. To maintain parsimony and interpretability, random effects were therefore restricted to the intercept and two dynamic covariates (relative speed and relative distance), while other vehicle dynamics and geometric features were retained as fixed effects, following recommendations from (Anis et al., 2025a). This choice reflects theoretical expectations: vehicle dynamics are highly variable across intersections depending on signal phasing, approach alignment, and driver behavior, whereas roadway structure tends to exert a more stable influence. The resulting specification captures both site-specific dynamics and broader geometric constraints, group-specific random effects illustrated in Fig. 5.1.1 with posterior summaries reported in Table 5.
(a) $\beta_{\mu,0[1,10]}$
(b) $\gamma_{\mu,1[1,10]}$
(c) $\gamma_{\mu,2[1,10]}$
(d) $\beta_{\sigma,0[1,10]}$
(e) $\gamma_{\sigma,1[1,10]}$
(f) $\gamma_{\sigma,2[1,10]}$
(g) $\beta_{\xi,0[1,10]}$
Fig. 9.: Posterior estimates for V-V near-misses at intersections
Table 5: Posterior estimates for V-V near misses at intersections
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -1.870 0.039 $1.917,-1.811]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.893 0.186 $2.309,-1.572]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative speed - - - 0.745 0.181 $0.393,1.134]^{\dagger}$
$\gamma_{\mu,2,k}$ Random Relative distance - - - -0.491 0.157 $0.819,-0.280]^{\dagger}$
$\beta_{\mu,1}$ Fixed Relative speed 0.639 0.050 $0.246,0.739]^{\dagger} - - -$
$\beta_{\mu,2}$ Fixed Relative acceleration 1.064 0.074 $0.126,2.018]^{\dagger} -$0.155 0.032 $0.221,-0.093]^{\dagger}$
$\beta_{\mu,3}$ Fixed Relative deceleration -0.013 0.037 [-0.088, 0.059] 0.140 0.063 $0.004,0.238]^{\dagger}$
$\beta_{\mu,4}$ Fixed Relative distance -0.442 0.036 $0.505,-0.418]^{\dagger} - - -$
$\beta_{\mu,5}$ Fixed Jerk -0.012 0.006 $021,-0.002]^{\dagger} -0.0$27 0.015 $0.046,-0.006]^{\dagger}$
$\beta_{\mu,6}$ Fixed Heading difference 0.024 0.026 [-0.029, 0.072] 0.028 0.039 [-0.052, 0.097]
$\beta_{\mu,7}$ Fixed Steering difference -0.004 0.049 [-0.096, 0.092] 0.059 0.026 $0.003,0.107]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume -0.003 0.005 [-0.012, 0.006] -0.018 0.035 [-0.085, 0.051]
$\beta_{\mu,9}$ Fixed Turn movement (Left =1, else=0) 0.086 0.026 [-0.102, 0.242] 0.073 0.011 $0.042,0.205]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\sigma,0}$ Fixed Intercept -0.571 0.058 $0.680,-0.455]^{\dagger} - - -$
$\beta_{\sigma,0,k}$ Random Intercept - - - -0.599 0.311 $1.240,-0.056]^{\dagger}$
$\gamma_{\sigma,1,k}$ Random Relative speed - - - -0.265 0.254 $0.810,-0.043]^{\dagger}$
$\gamma_{\sigma,2,k}$ Random Relative distance - - - -0.653 0.297 $1.121,-0.033]^{\dagger}$
$\beta_{\sigma,1}$ Fixed Relative speed 0.012 0.084 [-0.139, 0.183] - - -
$\beta_{\sigma,2}$ Fixed Relative acceleration -1.686 0.865 $3.066,-0.627]^{\dagger} -$0.226 0.066 $0.360,-0.108]^{\dagger}$
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.218 0.051 $0.321,-0.120]^{\dagger} -$0.253 0.053 $0.356,-0.151]^{\dagger}$
$\beta_{\sigma,4}$ Fixed Relative distance -0.062 0.008 $0.078,-0.046]^{\dagger} - - -$
$\beta_{\sigma,5}$ Fixed Jerk 0.019 0.009 $005,0.030]^{\dagger} 0.0$12 0.003 $0.007,0.018]^{\dagger}$
$\beta_{\sigma,6}$ Fixed Heading difference -0.063 0.071 [-0.212, 0.060] -0.055 0.014 $0.133,-0.026]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Steering difference 0.063 0.065 [-0.061, 0.191] -0.053 0.040 $0.117,-0.011]^{\dagger}$
$\beta_{\sigma,8}$ Fixed Volume -0.001 0.007 [-0.015, 0.013] -0.004 0.044 [-0.090, 0.084]
$\beta_{\sigma,9}$ Fixed Turn movement(Left =1, else=0) 0.042 0.120 [-0.183, 0.285] 0.217 0.123 $0.025,0.471]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.305 0.051 $0.413,-0.208]^{\dagger} - - -$
$\beta_{\xi,0,k}$ Random Intercept - - - -0.684 0.253 $1.309,-0.314]^{\dagger}$
Model Fit
DIC 526.9 508.9
WAIC 539.4 521.5
LOOIC 540.1 522.3
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
The fixed-parameter benchmark (HBSFP) provides a baseline against which the grouped random specification can be compared. Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, underscoring the explanatory value of allowing relative speed and spacing effects to vary across intersection clusters. The direction and magnitude of these coefficients provide valuable insights into the mechanisms underlying COR. The random-parameter specification highlights the spatially varying influence of relative speed and relative distance on both the location and scale components of the HBSGRP distribution. For $\mu$, relative speed shows a positive effect across intersections, indicating that larger speed differentials between interacting vehicles lead to sharply reduced 2D-TTC values and more severe near-miss conditions. In contrast, relative distance is negatively associated with $\mu$, meaning that shorter headways systematically increase near-miss severity by limiting the buffer available for evasive action. For $\sigma$, both covariates are associated with reduced variability in near-miss severity. Relative speed has a negative coefficient, suggesting that high-speed interactions produce consistently critical outcomes, dangerous but less variable, because drivers have little opportunity to recover through braking or swerving. Similarly, relative distance exerts a negative effect on $\sigma$, reflecting that short headways yield uniformly critical near-miss with little dispersion in outcomes. Together, these results confirm the physical intuition that speed amplifies severity while distance compresses variability, producing deterministic near miss outcomes under constrained conditions. Fig. 5.1.1 further illustrate this spatial heterogeneity across intersection, while most locations show consistent trends, the I2 site exhibits unusually wide CI for the $\mu$ effects and CI exceed zero for the $\sigma$, across intercept, relative speed, and relative distance. This instability is directly attributable to I2’s small BM sample size, which weakens posterior precision and introduces greater uncertainty in estimated group-level effects. Despite this, the overall pattern across intersections remains robust: relative speed consistently intensifies near-miss severity, while shorter distances exacerbate risk and constrain outcome variability.
Beyond the random effects, several covariates were modeled as fixed parameters due to their relatively stable contributions across intersections. Their interpretation reveals distinct roles of longitudinal dynamics, lateral control, maneuver type, and exposure effects in shaping near-miss severity and variability. Table 5 indicate that relative acceleration shows a significant negative effect on both the location and scale parameters. This suggests that acceleration is generally associated with smoother and less variable near-miss, consistent with situations where vehicles accelerate out of a green phase or clear an intersection without abrupt interaction. By contrast, relative deceleration exerts a positive effect on severity ($\mu$) but a negative effect on variability ($\sigma$). This indicates that abrupt braking sharply intensifies near-misses but also produces more uniform outcomes, reflecting the deterministic nature of emergency deceleration in high-risk encounters. Jerk, which captures the rate of change in acceleration, has the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$. Although jerk slightly reduces average near-miss severity (perhaps due to hesitation or driver corrections), it simultaneously increases unpredictability in outcomes, highlighting the unstable nature of rapid throttle/brake oscillations.
Lateral vehicle dynamics also contribute meaningfully to near-miss risk estimation. Steering difference emerges as a significant contributor in the HBSGRP specification, showing a positive association with $\mu$ and a negative association with $\sigma$. This suggests that abrupt steering maneuvers intensify near-miss severity but lead to more consistent outcomes within constrained intersection geometries. Heading difference also shows a significant positively association with $\mu$ and a negatively with $\sigma$. Indicating that even small misalignments in trajectory lead to predictable patterns of near-miss outcomes, consistent with the rigid geometry of signalized approaches.
The categorical indicator for left-turn movements is significant and has a positive effect on both $\mu$ and $\sigma$. This aligns with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies. The result suggests that left turns not only elevate near-miss severity but also introduce greater variability in outcomes, reflecting the wide range of intersection designs and driver behaviors under turning conditions. Traffic volume is statistically insignificant in both $\mu$ and $\sigma$. Nonetheless, it was retained in the model to ensure conceptual completeness, since it represents exposure. The lack of effect implies that short-duration 2D-TTC near-miss events are governed more by instantaneous vehicle interactions than by aggregate traffic demand.
Finally, the estimated shape parameter ($\xi$) is negative and significant across all intersections (Fig. 8(g)), indicating a bounded upper tail of the UGEV distribution. This supports the Weibull domain as the appropriate tail form for modeling extreme near-misses in signalized urban settings. Importantly, group-level variability in $\xi$ suggests that intersection design and control strategies influence not only the frequency but also the extremal nature of near-miss outcomes.
The HBSGRP model for V–I near-misses specifies the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in near-miss severity and variability across intersections. Relative speed (location) and relative distance (scale) were also treated as random effects (Fig. 5.1.1), though their 95% CI overlapped zero. They were retained on theoretical grounds, given their inclusion improve overall metrics. All other covariates were modeled as fixed effects, representing more stable behavioral and geometric influences. Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC over the HBSFP benchmark, suggesting that the grouped random parameter structure adds nuance but yields smaller performance gains compared to the V–V case.
Table 6: Posterior estimates for V-I near misses at intersections
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -1.392 0.039 $1.461,-1.299]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.401 0.092 $1.598,-1.236]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative speed - - - 0.025 0.020 [-0.012, 0.065]
$\beta_{\mu,1}$ Fixed Relative speed -0.012 0.022 [-0.054, 0.031] - - -
$\beta_{\mu,2}$ Fixed Relative acceleration -0.020 0.023 [-0.067, 0.022] -0.020 0.024 [-0.068, 0.029]
$\beta_{\mu,3}$ Fixed Relative deceleration -0.066 0.013 $075,-0.055]^{\dagger} -0.0$26 0.011 $0.071,-0.009]^{\dagger}$
$\beta_{\mu,4}$ Fixed Relative distance -0.357 0.035 $0.427,-0.290]^{\dagger} -$0.359 0.026 $0.405,-0.306]^{\dagger}$
$\beta_{\mu,5}$ Fixed Jerk 0.270 0.355 [-0.457, 0.556] -0.021 0.005 [-0.026, 0.01]
$\beta_{\mu,6}$ Fixed Heading difference 0.040 0.022 $005,0.084]^{\dagger} 0.0$62 0.026 $0.010,0.127]^{\dagger}$
$\beta_{\mu,7}$ Fixed Steering difference 0.009 0.018 [-0.027, 0.045] 0.038 0.022 $0.014,0.074]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume -0.017 0.021 [-0.061, 0.023] -0.023 0.023 [-0.069, 0.027]
$\beta_{\mu,9}$ Fixed Turn movement (Left =1, else=0) -0.129 0.080 $0.265,-0.061]^{\dagger} -$0.579 0.310 $1.231,-0.220]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\sigma,0}$ Fixed Intercept -0.100 0.038 $0.167,-0.009]^{\dagger} - - -$
$\beta_{\sigma,0,k}$ Random Intercept - - - -0.197 0.098 $0.374,-0.001]^{\dagger}$
$\gamma_{\sigma,1,k}$ Random Relative distance - - - 0.042 0.034 [-0.028, 0.107]
$\beta_{\sigma,1}$ Fixed Relative speed 0.036 0.019 $013,0.063]^{\dagger} 0.0$40 0.019 $0.001,0.076]^{\dagger}$
$\beta_{\sigma,2}$ Fixed Relative acceleration 0.012 0.015 [-0.015, 0.042] 0.046 0.011 $0.009,0.084]^{\dagger}$
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.147 0.056 $0.285,-0.027]^{\dagger}$g -0.177 0.056 $0.289,-0.047]^{\dagger}$
$\beta_{\sigma,4}$ Fixed Relative distance 0.209 0.038 $0.140,0.282]^{\dagger} - - -$
$\beta_{\sigma,5}$ Fixed Jerk 0.611 0.133 $0.227,0.886]^{\dagger}$ 0.109 0.033 $0.016,0.156]^{\dagger}$
$\beta_{\sigma,6}$ Fixed Heading difference -0.023 0.018 $057,-0.004]^{\dagger} -0.0$90 0.021 $0.130,-0.049]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Steering difference -0.012 0.018 [-0.045, 0.023] -0.024 0.020 $0.062,-0.017]^{\dagger}$
$\beta_{\sigma,8}$ Fixed Volume 0.004 0.016 [-0.028, 0.034] 0.029 0.018 $0.003,0.064]^{\dagger}$
$\beta_{\sigma,9}$ Fixed Turn movement (Left =1, else=0) 0.100 0.053 $0.019,0.197]^{\dagger}$ 0.264 0.049 $0.163,0.349]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.677 0.042 $0.749,-0.594]^{\dagger} - - -$
$\beta_{\xi,0,k}$ Random Intercept - - - -0.680 0.096 $0.876,-0.501]^{\dagger}$
Model Fit
DIC 3197 3179.6
WAIC 3202 3194.4
LOOIC 3212.87 3209.8
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
(a) $\beta_{\mu,0[1,10]}$
(b) $\beta_{\sigma,0[1,10]}$
(c) $\beta_{\xi,0[1,10]}$
Fig. 10.: Posterior estimates for V-I near-misses at intersections
In the location parameter, several interpretable effects emerge. Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC, elevating severity. Relative deceleration also carries a significant negative sign, which contrasts with its positive effect in V–V near-miss. This suggests that braking near infrastructure reflects proactive yielding or evasive action, thereby limiting severity rather than escalating it. Heading and steering differences are positive and significant, indicating that lateral deviations (e.g., turning toward or away from curbs) increase near-miss severity. Left-turn movement, however, shows a significant negative association, suggesting that left-turn trajectories may provide greater clearance to boundaries, reducing average severity compared to right turns. This finding requires careful interpretation, as it may partly reflect geometric factors such as wider turning radii or reduced pedestrian/boundary overlap in left-turn phasing. Relative speed, though included as a random effect, is not significant in the pooled analysis, highlighting that V–I risks are less sensitive to speed differentials than V–V encounters.
The scale parameter results add further insight into variability. The negative random intercept suggests baseline differences across intersections, with some sites producing more predictable outcomes than others. Relative speed exhibits a positive and significant association with scale, implying that faster approaches to infrastructure yield greater variability in severity—contrasting with the V–V case, where speed reduced dispersion. This difference reflects the inherently diverse driver responses to boundary proximity: some brake aggressively, while others take wider trajectories or maintain speed. Relative acceleration is also positive and significant, reinforcing the idea that heterogeneous driver behavior during acceleration near boundaries widens outcome variability. In contrast, relative deceleration has a significant negative effect on scale, meaning that strong braking compresses the severity distribution into consistently critical outcomes. Jerk is similarly positive, confirming that abrupt longitudinal control amplifies variability. Lateral dynamics again exhibit distinct roles: heading and steering differences reduce variability, suggesting that once a driver commits to a lateral maneuver, outcomes become more predictable. Finally, left-turn movement increases variability despite lowering average severity, consistent with heterogeneous geometric layouts and driver strategies in left-turn contexts.
The shape parameter remains negative and significant across intersections, supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail. This implies a physical or behavioral limit to how severe boundary near-misses can become, constrained by roadway geometry and driver response capacity. Spatial variation in the shape parameter underscores the role of intersection design and infrastructure placement in shaping the extremal properties of near-miss distributions. Overall, the intersection-level results suggest that V–I near-misses are driven more by clearance, braking behavior, and lateral maneuvers than by speed differentials. Unlike V–V near-misses, where high speeds produce deterministic severity, V–I cnear-misses are marked by greater variability, reflecting the complex interplay between infrastructure geometry, driver avoidance strategies, and intersection control.
5.2 Segment-level
The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical framework (Table 7). Allowing the intercept and three time-varying vehicle-based measures, relative speed, relative deceleration, and relative distance, to vary across segments (See Fig. 5.2). This specification acknowledges that longitudinal dynamics in uninterrupted midblock contexts are shaped not only by vehicle behavior but also by geometric features such as lane configuration, median presence, and driveway density. All remaining covariates, including relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density, were modeled as fixed effects due to their comparatively stable influence across segments. The HBSGRP specification produces substantial improvements in model fit over the HBSFP baseline, confirming the value of partial pooling in capturing heterogeneity across the nine directional segments analyzed.
For the location parameter ($\mu$), random effects highlight systematic heterogeneity across the nine directional segments. Relative speed exerts a significant positive effect, confirming that higher closing speeds consistently compress available 2D-TTC and elevate near-miss severity. However, the posterior distributions in Fig. 5.2 reveal that the magnitude of this effect varies substantially across segments. Segments with uninterrupted, high-speed flows (e.g., S11) show stronger effects, while others with denser access points (e.g., S6, S7 and S13) exhibit wider credible intervals. Relative deceleration also carries a negative sign, though with more overlap in its credible intervals across segments, suggesting that emergency braking may sometimes mitigate severity in smoother-flow environments, while in higher-speed segments it intensifies near-misses. Notably, in some lower-sample segments, credible intervals for the location intercept approach or cross zero, reflecting weaker estimation stability under sparse BM events. Relative distance displays a consistently negative effect across segments, meaning that shorter headways reliably reduce 2D-TTC, yet its posterior variation underscores differences in spacing behavior along divided versus undivided facilities.
Refer to caption Refer to caption
(a) $\beta_{\mu,0[1,18]}$ (b) $\gamma_{\mu,1[1,18]}$
Refer to caption Refer to caption
(c) $\gamma_{\mu,2[1,18]}$ (d) $\gamma_{\mu,3[1,18]}$
Refer to caption Refer to caption
(e) $\beta_{\sigma,0[1,18]}$ (f) $\gamma_{\sigma,1[1,18]}$
Refer to caption Refer to caption
(g) $\gamma_{\sigma,2[1,18]}$ (h) $\gamma_{\sigma,3[1,18]}$
(i) $\beta_{\xi,0[1,18]}$
Fig. 11.: Posterior estimates for V-V near-misses at segments
Table 7: Posterior estimates V-V near misses at segments
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -2.117 0.032 $2.181,-2.025]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.047 0.171 $1.369,-0.772]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative speed - - - 0.406 0.068 $0.286,0.543]^{\dagger}$
$\gamma_{\mu,2,k}$ Random Relative deceleration - - - -0.082 0.016 $0.197,-0.057]^{\dagger}$
$\gamma_{\mu,3,k}$ Random Relative distance - - - -0.600 0.082 $0.779,-0.454]^{\dagger}$
$\beta_{\mu,1}$ Fixed Relative speed 0.446 0.037 $0.376,0.519]^{\dagger} - - -$
$\beta_{\mu,2}$ Fixed Relative acceleration -0.068 0.017 $154,-0.034]^{\dagger} -0.0$79 0.037 $0.163,-0.032]^{\dagger}$
$\beta_{\mu,3}$ Fixed Relative deceleration -0.074 0.018 $0.108,-0.040]^{\dagger} - - -$
$\beta_{\mu,4}$ Fixed Relative distance -0.669 0.063 $0.801,-0.552]^{\dagger} - - -$
$\beta_{\mu,5}$ Fixed Jerk -0.004 0.001 $0.007,-0.002]^{\dagger} -$0.156 0.043 $0.234,-0.101]^{\dagger}$
$\beta_{\mu,6}$ Fixed Heading difference 0.043 0.016 $011,0.073]^{\dagger} 0.0$44 0.008 $0.029,0.059]^{\dagger}$
$\beta_{\mu,7}$ Fixed Steering difference 0.042 0.012 $018,0.066]^{\dagger} 0.0$26 0.010 $0.007,0.046]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume 0.035 0.013 $009,0.060]^{\dagger} 0.0$19 0.008 $0.002,0.038]^{\dagger}$
$\beta_{\mu,9}$ Fixed Lane no -0.006 0.011 [-0.027, 0.015] 0.009 0.011 [-0.011, 0.032]
$\beta_{\mu,10}$ Fixed Lane width 0.003 0.001 $001,0.007]^{\dagger} 0.0$17 0.009 $0.002,0.031]^{\dagger}$
$\beta_{\mu,11}$ Fixed Driveway density -0.001 0.015 [-0.029, 0.027] 0.489 0.114 $0.256,0.578]^{\dagger}$
$\beta_{\mu,12}$ Fixed Median (Undivided=1, else=0) -0.004 0.033 [-0.070, 0.062] -0.412 0.155 $0.711,-0.070]^{\dagger}$
$\beta_{\mu,13}$ Fixed Vehicle position (Left lane=1, else=0) 0.029 0.038 [-0.048, 0.103] 0.055 0.006 $0.012,0.105]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -0.622 0.045 $0.711,-0.533]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -0.253 0.047 $0.952,-0.156]^{\dagger}$
$\gamma_{\sigma,1,k}$ Random Relative speed - - - -0.055 0.026 $0.107,-0.003]^{\dagger}$
$\gamma_{\sigma,2,k}$ Random Relative deceleration - - - -0.119 0.018 $0.154,-0.084]^{\dagger}$
$\gamma_{\sigma,3,k}$ Random Relative distance - - - -0.304 0.117 $0.520,-0.084]^{\dagger}$
$\beta_{\sigma,1}$ Fixed Relative speed -0.116 0.050 $0.217,-0.023]^{\dagger} - - -$
$\beta_{\sigma,2}$ Fixed Relative acceleration -0.611 0.097 $817,-0.439]^{\dagger} -0.0$24 0.173 [-0.277, 0.298]
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.060 0.019 $0.097,-0.024]^{\dagger} - - -$
$\beta_{\sigma,4}$ Fixed Relative distance 0.003 0.064 [-0.121, 0.130] - - -
$\beta_{\sigma,5}$ Fixed Jerk 0.006 0.000 $0.004,0.008]^{\dagger}$ 0.357 0.147 $0.061,0.611]^{\dagger}$
$\beta_{\sigma,6}$ Fixed Heading difference 0.065 0.017 $2,0.099]^{\dagger} 0.033 0.0$11 $0.017,0.068]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Steering difference 0.011 0.016 [-0.020, 0.043] -0.001 0.019 [-0.031, 0.029]
$\beta_{\sigma,8}$ Fixed Volume -0.011 0.016 [-0.043, 0.020] -0.012 0.015 [-0.030, 0.029]
$\beta_{\sigma,9}$ Fixed Lane no -0.031 0.014 $049,-0.005]^{\dagger} -0.0$28 0.017 $0.061,-0.002]^{\dagger}$
$\beta_{\sigma,10}$ Fixed Lane width -0.006 0.003 $012,-0.000]^{\dagger} -0.0$15 0.009 $0.044,-0.001]^{\dagger}$
$\beta_{\sigma,11}$ Fixed Driveway density 0.040 0.018 $0.004,0.075]^{\dagger}$ 0.284 0.215 $0.019,0.648]^{\dagger}$
$\beta_{\sigma,12}$ Fixed Median (Undivided=1, else=0) 0.076 0.042 [-0.006, 0.158] -0.205 0.095 $0.369,-0.058]^{\dagger}$
$\beta_{\sigma,13}$ Fixed Vehicle position (Left lane=1, else=0) -0.049 0.052 [-0.156, 0.051] -0.109 0.052 $0.209,-0.006]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.116 0.030 $0.178,-0.059]^{\dagger} - - -$
$\beta_{\xi,0,k}$ Random Intercept - - - -0.185 0.144 $0.436,-0.163]^{\dagger}$
Model Fit
DIC 4528 4189
WAIC 4594 4276
LOOIC 4608.25 4278
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
Among the fixed effects, relative acceleration shows a significant negative coefficient in both models, implying that acceleration phases are systematically linked to higher near-miss severity on segments. This contrasts with intersection contexts, where acceleration often reflected flow resumption. Here, aggressive acceleration during overtaking or merging maneuvers reduces spacing and increases risk. Jerk is also negative, reinforcing the role of abrupt longitudinal control in producing critical outcomes. Lateral control indicators—heading and steering differences—show positive signs, suggesting that lateral adjustments modestly increase 2D-TTC where lane-changing flexibility exists. Driveway density exhibits a strong positive effect on $\mu$, which may appear counterintuitive, but is likely explained by anticipatory driver caution in access-rich environments. Conversely, undivided medians sharply reduce $\mu$, indicating that exposure to opposing flows creates consistently severe interactions.
For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, showing that extreme dynamics compress outcome variability, leading to uniformly critical near misses. The plots confirm this: high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, with narrow posteriors in segments such as S12 and S15. In contrast, segments like S10 display wider CI for random scale effects, again reflecting limited BM sample sizes. Fixed effects add further nuance: jerk and heading difference increase scale, suggesting that volatile acceleration or lateral corrections introduce heterogeneity in outcomes, while lane number and lane width reduce scale, indicating that greater maneuvering space fosters more predictable near-misses. Driveway density increases dispersion, consistent with the variability introduced by frequent merging and crossing maneuvers. Median type again exerts a strong negative influence, reinforcing that the absence of raised medians compresses severity outcomes into consistently critical ranges.
Finally, the shape parameter ($\xi$) remains negative and significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain. This indicates physical and behavioral limits to maximum 2D-TTC values during critical segment near-misses, constrained by speed environment, geometry, and driver reaction times. Posterior variation in the shape intercept suggests that undivided segments experience steeper extremal limits, while divided facilities allow marginally greater spread in outcomes. Overall, the segment-level results highlight the primacy of longitudinal dynamics in shaping corridor risk, with acceleration, braking, and headway dominating near-miss severity, while roadway design elements such as medians, lane width, and access density govern both severity and variability. The superior fit of the HBSGRP model underscores the importance of capturing these segment-specific dynamics, as fixed-parameter structures fail to account for the spatially contingent influence of speed and spacing across midblock contexts.
The HBSGRP specification for segment-level V–I near-misses provides a detailed view of how longitudinal dynamics and roadway design influence the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig. 5.2). In this model, the intercept and relative distance are treated as group-level random effects in both the location and scale components of the HBSGRP model, allowing for spatial heterogeneity in how spacing between interacting vehicles shapes COR across segments. All other predictors—including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence, are modeled as fixed effects, reflecting their expected stability across segment environments.
(a) $\beta_{\mu,0[1,18]}$
(b) $\gamma_{\mu,1[1,18]}$
(c) $\beta_{\sigma,0[1,18]}$
(d) $\gamma_{\sigma,1[1,18]}$
(e) $\beta_{\xi,0[1,18]}$
Fig. 12.: Posterior estimates V-I near misses at segments
Table 8: Posterior estimates V-I near misses at segments
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -1.287 0.037 $1.352,-1.22]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.263 0.122 $1.494,-1.037]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative distance - - - -0.317 0.148 $0.632,-0.104]^{\dagger}$
$\beta_{\mu,1}$ Fixed Relative speed -0.027 0.018 $061,-0.008]^{\dagger} 0.0$09 0.001 $0.006,0.027]^{\dagger}$
$\beta_{\mu,2}$ Fixed Relative acceleration 8.36E-04 0.019 [-0.045, 0.036] 0.025 0.013 $0.002,0.057]^{\dagger}$
$\beta_{\mu,3}$ Fixed Relative deceleration -0.039 0.008 $098,-0.004]^{\dagger} -0.0$32 0.008 $0.087,-0.014]^{\dagger}$
$\beta_{\mu,4}$ Fixed Relative distance -0.257 0.025 $0.304,-0.207]^{\dagger} - - -$
$\beta_{\mu,5}$ Fixed Jerk -0.124 0.069 [-0.224, 0.289 ] -0.074 0.039 [-0.124, 0.202]
$\beta_{\mu,6}$ Fixed Heading difference 0.025 0.025 [-0.025, 0.069] 0.027 0.014 $0.001,0.049]^{\dagger}$
$\beta_{\mu,7}$ Fixed Steering difference 0.033 0.017 $004,0.065]^{\dagger} 0.0$41 0.016 $0.003,0.066]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume -0.002 0.007 [-0.016,0.010] 0.043 0.021 $0.015,0.084]^{\dagger}$
$\beta_{\mu,9}$ Fixed Lane no -0.001 0.012 [-0.019, 0.016] 0.042 0.021 $0.004,0.078]^{\dagger}$
$\beta_{\mu,10}$ Fixed Lane width -0.027 0.018 $061,-0.008]^{\dagger} 0.0$13 0.009 $0.000,0.032]^{\dagger}$
$\beta_{\mu,11}$ Fixed Driveway density -0.007 0.024 [-0.057, 0.042] -0.237 0.105 $0.403,-0.020]^{\dagger}$
$\beta_{\mu,12}$ Fixed Median (Undivided=1, else=0) -0.644 0.102 $773,-0.513]^{\dagger} -0.0$83 0.059 $0.196,-0.019]^{\dagger}$
$\beta_{\mu,13}$ Fixed Vehicle position (Left lane=1, else=0) 0.057 0.033 $028,0.091]^{\dagger} -0.0$64 0.014 $0.082,-0.038]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -0.495 0.156 $0.703,-0.303]^{\dagger} - - -$
$\beta_{\sigma,0,k}$ Random Intercept - - - -0.076 0.104 [-0.265, 0.083]
$\gamma_{\sigma,1,k}$ Random Relative distance - - - 0.135 0.103 $0.011,0.318]^{\dagger}$
$\beta_{\sigma,1}$ Fixed Relative speed 0.039 0.016 $012,0.061]^{\dagger} 0.0$13 0.001 $0.004,0.038]^{\dagger}$
$\beta_{\sigma,2}$ Fixed Relative acceleration 0.002 0.017 [-0.030, 0.042] -0.026 0.014 $0.064,-0.001]^{\dagger}$
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.036 0.013 $075,-0.014]^{\dagger} -0.0$48 0.021 $0.089,-0.022]^{\dagger}$
$\beta_{\sigma,4}$ Fixed Relative distance 0.167 0.018 $0.136,0.199]^{\dagger} - - -$
$\beta_{\sigma,5}$ Fixed Jerk -0.108 0.054 [0.452, -0.119] -0.078 0.119 [-0.312, 0.078]
$\beta_{\sigma,6}$ Fixed Heading difference -0.019 0.020 [-0.053, 0.020] -0.029 0.017 $0.049,-0.001]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Realative steering angle -0.033 0.012 $065,-0.001]^{\dagger} -0.0$32 0.024 $0.059,-0.003]^{\dagger}$
$\beta_{\sigma,8}$ Fixed Volume -0.007 0.018 [-0.039, 0.027] 0.001 0.016 [-0.026, 0.030]
$\beta_{\sigma,9}$ Fixed Lane no -0.028 0.017 $058,-0.006]^{\dagger} -0.0$21 0.012 $0.046,-0.002]^{\dagger}$
$\beta_{\sigma,10}$ Fixed Lane width -0.010 0.020 [-0.044, 0.030] 0.018 0.016 $0.010,0.045]^{\dagger}$
$\beta_{\sigma,11}$ Fixed Driveway density 0.016 0.019 [-0.021, 0.055] -0.125 0.028 [-0.239, 0.018]
$\beta_{\sigma,12}$ Fixed Median (Undivided=1, else=0) 0.663 0.166 $0.440,0.880]^{\dagger}$ 0.174 0.090 $0.263,0.553]^{\dagger}$
$\beta_{\sigma,13}$ Fixed Vehicle position (Left lane=1, else=0) -0.046 0.027 $080,-0.012]^{\dagger} 0.0$55 0.010 $0.039,0.071]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.845 0.028 $0.897,-0.788]^{\dagger}$g - - -
$\beta_{\xi,0,k}$ Random Intercept - - - -0.887 0.118 $1.123,-0.659]^{\dagger}$
Model Fit
DIC 6760 6550
WAIC 6775 6595
LOOIC 6791.2 6601
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
For the location parameter, shorter relative distances strongly reduce 2D-TTC values, confirming that closer spacing consistently produces more critical near-miss outcomes. The random effect for distance further indicates that influence varies by segment group, suggesting sensitivity to local geometric or operational conditions. Relative speed displays a notable divergence: while the fixed-parameter benchmark shows a negative coefficient, the grouped model produces a small but positive estimate, implying that in certain contexts, higher speeds are associated with marginally longer 2D-TTC values. This reversal highlights that segment-level heterogeneity in speed environments, such as free-flow arterials versus constrained multilane facilities, can fundamentally alter the role of speed in shaping severity. Relative acceleration emerges as a significant positive effect in the grouped specification, suggesting that acceleration phases along segments may often represent clearance or recovery maneuvers rather than heightened near-miss risk. In contrast, relative deceleration maintains a negative coefficient across both models, reflecting the reality that braking during close-following conditions is tied to critical reductions in 2D-TTC, likely triggered by sudden evasive actions.
Several geometric and operational features are also important for location. Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments yield small gains in clearance and mitigate near-miss severity. Volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that higher-capacity, wider facilities afford drivers greater maneuvering space, leading to slightly larger 2D-TTC values. The effect of driveway density, however, is particularly striking: while insignificant in the fixed-parameter model, the grouped specification produces a large negative estimate with a wide CI. This instability suggests that driveway effects are highly localized—segments with dense access points may concentrate more severe near-miss outcomes, but the data are sparse and sensitive to segment-specific characteristics. Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing traffic flows. The left-lane variable shows a sign reversal, with the fixed model suggesting higher severity in interior lanes and the grouped model indicating greater severity in outer lanes. This again illustrates heterogeneity across segment groups, where the positional context of a lane relative to roadside boundaries and merging activity shifts the locus of risk.
The scale parameter provides complementary insights into the dispersion of near-miss severity. The random effect for relative distance is positive, showing that tighter spacing increases not only severity but also variability in outcomes, as different drivers adopt divergent strategies under minimal headway conditions. Relative speed also has a positive association with scale, indicating that higher speeds amplify heterogeneity in outcomes across V–I interactions, unlike in V–V segment near-misses, where speed reduces dispersion. Acceleration, by contrast, has a negative coefficient in the grouped model, suggesting that acceleration phases yield more predictable outcomes, consistent with smoother-flow recovery dynamics. Relative deceleration also reduces variability, implying that braking in near-crash contexts converges toward uniformly critical outcomes.
Geometric attributes reveal nuanced effects on variability. Lane width exerts a positive effect, suggesting that while wider lanes modestly reduce average severity, they also introduce greater variability in outcomes by enabling diverse driver responses. The undivided median type strongly increases scale, underscoring that undivided facilities not only heighten the average risk but also broaden the range of severity outcomes, likely due to variation in driver avoidance maneuvers. The left-lane effect is positive in scale, indicating that severity outcomes in these lanes are more heterogeneous, perhaps reflecting the diversity of driver strategies in navigating outer versus inner lanes. Driveway density, although statistically unstable, shows a wide and imprecise effect that suggests variability is strongly contingent on site-specific access management patterns.
The shape parameter remains negative and significant across both specifications, confirming that the upper tail of the 2D-TTC distribution is bounded. This supports the Weibull characterization of extremes, implying that even in segment-level V–I interactions, physical and behavioral limits constrain how severe the most critical near-miss events can be. Variation in shape across segment groups highlights the role of local geometry, access density, and operational features in shaping the extremal properties of near-miss distributions.
Together, the segment-level analyses of V–V and V–I near misses underscore that roadway geometry and vehicle dynamics interact to shape both the severity and variability of extreme near-misses. For V–V events, relative speed and distance dominate as key random effects, with speed consistently amplifying severity and distance compressing variability, reflecting deterministic outcomes under close-following conditions. For V–I events, however, the role of speed is more heterogeneous—sometimes protective, sometimes risk-enhancing—depending on the segment context. Access-related variables, particularly driveway density and median type, emerge as critical for V–I outcomes, underscoring the importance of access management and median design in shaping roadside safety. In both near-miss types, the negative and significant shape parameter indicates bounded extremes, reinforcing the physical and behavioral limits to crash risk escalation.
6 Model validation
The dataset for the study corridor provides precise geolocation for each extreme near-miss event and readily available annual corridor-level observed crashes. The dataset for the study corridor provides precise geolocation for each extreme near-miss event from Eqn. 32. To address this limitation, validation was conducted as an event-level case–control approach, designed to assess whether the model assigns higher predicted probabilities to more severe near-miss events. Posterior exceedance probabilities were computed from Eqn. 30 for severity thresholds ranging from –0.1 to –0.9 seconds, and these probabilities were then used to evaluate classification performance using the area under the receiver operating characteristic curve (ROC–AUC). This metric is particularly well-suited to rare-event prediction because it is independent of a specific threshold.
(a) V–V Segment
(b) V–I Segment
(c) V–V Intersection
(d) V–I Intersection
Fig. 13.: ROC curves for four HBSGRP–UGEV models across severity thresholds from $\omega=-0.1$s to $\omega=-0.9$s
ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V, V–I interactions at segment and intersection levels. The results (Figure 6) reveal clear distinctions across near-miss types and roadway contexts. The segment-level V-V model consistently achieved the highest accuracy, with mean AUC values of about 0.89 and very little variation across thresholds. This indicates stable and robust predictive performance, reflecting the ability of grouped random parameters to capture high-speed longitudinal near-miss dynamics along uninterrupted midblock segments. The V-I model at the segment level achieved slightly lower accuracy, averaging around 0.79, and exhibited moderate sensitivity to the choice of severity threshold, a finding consistent with the greater variability in driver responses when vehicles approach roadside boundaries. At intersections, the V-V model demonstrated moderately strong predictive accuracy, with an average AUC of nearly 0.82 and relatively low volatility across thresholds. This indicates that grouped speed and spacing dynamics remain reliable even under more complex signalized and turning conditions. By contrast, the intersection-level V-I model performed weakest, averaging an AUC of approximately 0.75 and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
Taken together, these findings confirm that the HBSGRP–UGEV framework provides strong and reliable performance for V-V interactions, particularly in segment environments, while predictions for V-I interactions are more sensitive to threshold selection and local variability. This suggests that segment-level V-V estimates are the most operationally reliable for proactive safety monitoring, whereas intersection-level V-I near-misses may require further refinement. Incorporating more detailed information on curb geometry, pedestrian exposure, and high-resolution vehicle–infrastructure proximity measures could improve stability and predictive accuracy in future applications.
7 Summary and conclusions
This study addresses the persistent challenge of accurately estimating short-term crash occurrence risk (COR) in complex, mixed urban traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability. Existing approaches are limited by oversimplified near-miss representations, a predominant focus on vehicle–vehicle (V–V) interactions, and insufficient integration of spatial heterogeneity in statistical modeling. To overcome these gaps, we propose a high-fidelity, geometry-aware, and statistically robust framework for real-time COR estimation that introduces a novel two-dimensional time-to-collision (2D-TTC) indicator. This metric accounts for vehicle dimensions, dynamic states, and precise geometric positioning, enabling accurate detection of extreme near-miss events from both V–V and vehicle–infrastructure (V–I) interactions.
Extreme events are translated into probabilistic risk estimates through a non-stationary Univariate Generalized Extreme Value (UGEV) framework implemented within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) structure. A corridor-based grouping strategy was employed to address event sparsity, pooling near-miss samples across intersections and directional segments while preserving local heterogeneity. Execution of the framework involved extracting high-resolution vehicle trajectories from the Argoverse-2 motion forecasting dataset, integrating them with HD maps and supplementary roadway information, and identifying near-miss events. The model incorporated both vehicle dynamics covariate (relative speed, acceleration, deceleration, jerk, heading difference, steering difference, and traffic volume) and geometric or operational covariates (lane number, lane width, driveway density, median type, turning movement, and vehicle lane position). Grouped random parameters were specified for temporally varying dynamics, relative speed, relative deceleration, and relative distance, to capture their context-dependent effects. For benchmarking, a baseline hierarchical Bayesian fixed-parameter (HBSFP) model was also estimated.
Results demonstrate that HBSGRP consistently outperformed the HBSFP across both intersections and segments, achieving lower DIC, WAIC, and LOOIC values and providing greater estimation stability. Relative speed emerged as the dominant predictor of near-miss severity in intersections, while relative distance was more influential for segment-level V–V events. Deceleration played a critical role in segment-level V–V near-miss, underscoring the importance of longitudinal control. Fixed covariates such as heading difference, steering difference, lane width, driveway density, and median type also showed significant effects, demonstrating the ability of the model to capture both dynamic behaviors and roadway design influences.
Validation results further confirmed model credibility. ROC–AUC analysis showed strong predictive performance, with V–V segment models achieving the highest accuracy (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82). V–I models performed reasonably well (segment $\approx$ 0.79; intersection $\approx$ 0.75) but displayed greater sensitivity to severity threshold choice. These findings highlight that while V–V interactions are modeled with high stability, V–I near-misses especially at intersections, remain more complex and variable.
Despite its contributions, the study has several limitations. First, potential bias may arise from the dependence on AV onboard sensors, which may affect traffic state accuracy and near-miss detection. Second, the 2D-TTC formulation, while tractable, simplifies nonlinear vehicle dynamics, which could understate risks in extreme evasive maneuvers. Third, the current framework does not explicitly model near-miss severity levels. Future research should extend the framework to multivariate UGEV models incorporating complementary surrogate measures such as post-collision velocity change ($\Delta V$) and explicitly address vulnerable road user (VRU) near-misses. Broader validation across multiple cities, transferability calibration, and comparison with peaks-over-threshold (POT) GPD methods will further strengthen robustness and generalizability. As autonomous vehicle deployment increases, access to larger and richer trajectory datasets will enhance the statistical power of extreme value analysis and enable more nuanced safety assessments. In practice, the proposed HBSGRP–UGEV framework offers a scalable, data-driven tool for proactive safety monitoring at the corridor level. By enabling early identification of high-risk conditions and supporting real-time traffic management, it provides a pathway toward Vision Zero objectives, reducing fatalities and serious injuries while promoting safe, efficient, and equitable mobility in urban environments.
Acknowledgement:
This research is funded by Federal Highway Administration (FHWA) Exploratory Advanced Research 693JJ323C000010. The results do not reflect FHWA’s opinions.
Disclaimer:
The results presented in this document do not necessarily reflect those from the Federal Highway Administration.
Data availability:
Data and code will be made available on reasonable request.
[…]